Use the information in the table below:

Year Nominal GDP Real GDP

2012 280 290

2013 315 260

2014 305 310

  1. The increase in nominal GDP between 2012 and 2014 is:

1.1. -3.5 %

1.2. 19.2 %

1.3. 8.93 %

1.4. 6.90 %


2.The increase in real GDP between 2012 and 2014 is:

2.1. 5 %

2.2. 6.90 %

2.3. 7.14 %

2.4. 20 %




1
Expert's answer
2021-09-01T11:52:20-0400

Nominal GDP is the market value of goods and services produced in an economy, unadjusted for inflation. Real GDP is nominal GDP, adjusted for inflation to reflect changes in real output. Trends in the GDP deflator are similar to changes in the Consumer Price Index, which is a different way of measuring inflation.

Nominal GDP = Real GDP x GDP Deflator/100

In this case, the GDP Deflator (between 2012 and 2014 = Real GDP/Nominal GDP x 100

Nominal = 305 - 280 = 25

Real GDP = 301 - 290 = 11


11/25x 100 = 44

GDP Deflator = 44